A plane can fly 270 miles downwind in the same amount of time as it can travel 180 miles upwind. Find the velocity of the wind if the plane can fly 225 mph in still air
Answer by ptaylor(2198)

A plane can fly 270 miles downwind in the same amount of time as it can travel 180 miles upwind. Find the velocity of the wind if the plane can fly 225 mph in still air
Distance(d) equals Rate(r) times Time(t) or d=rt; r=d/t and t=d/r
Let v=velocity of wind
rate of plane downwind=225+v
rate of plane upwind=225-v
(t=d/r)
time travelled downwind=270/(225+v)
time travelled upwind=180/(225-v)
Now we are told that these times are equal
270/(225+v)=180/(225-v) cross multiply
270(225-v)=180(225+v) divide each side by 90
3(225-v)=2(225+v)
675-3v=450+2v
5v=225
v=45mph----velocity of wind
Hope this helps---ptaylor